OpenStudy (anonymous):

ordered pairs system of equations 4x+4=2 x-y=2

OpenStudy (anonymous):

like with the parenthesis.. and show me how u guys did it please {4x+4=2 {x-y=2

myininaya (myininaya):

so multiply bottom equation by 4

myininaya (myininaya):

now add the equations

myininaya (myininaya):

4x+4y=2 <---assume you meant this 4x-4y=8

myininaya (myininaya):

now add the equations

OpenStudy (anonymous):

yes... ty

myininaya (myininaya):

8x+0=10 8x=10 x=10/8 you can reduce and then just plug this value in the one of the equations given and solve for y

OpenStudy (anonymous):

u saved my life :-) i had a make up test on thisthat i didnt do on october

myininaya (myininaya):

:)

myininaya (myininaya):

wow so you understood everything i did?

OpenStudy (anonymous):

yeah.. its cuz i forgot

myininaya (myininaya):

great i'm glad i could help have a great night

OpenStudy (anonymous):

ty